# MMInA: Benchmarking Multihop Multimodal Internet Agents

### v1

@@ 4 Experiments
Model Agent Inputs Hop Success Rate ($\uparrow$) Task Success Rate ($\uparrow$)
@@
Caption CodeLLaMA-7B w/ caps 5.71 0 0 1.61 5.71 0 0 2.79
WebShop w/ caps Acc. Tree 29.72 0.00 0.00 5.61 29.72 0 0 14.55
Gemini-Pro w/ caps + Caps 30.12 11.09 0.05 12.38 30.12 1.52 0.38 15.22
GPT-4 w/ caps 38.58 20.70 3.43 13.50 38.58 3.79 0 19.85
Multimodal Fuyu-8B Images 27.36 0 0 5.52 27.36 0 0 13.39
Gemini-Pro-Vision + Acc. Tree 28.94 16.38 4.03 10.66 28.94 1.51 1.13 18.40
@@
Multimodal Gemini-Pro-Vision Images 39.17 23.93 4.78 14.27 39.17 10.61 1.13 20.13
(Memory) + Acc. Tree
Human - Webpage 99.02 97.91 93.77 98.43 99.02 95.34 88.12 96.25
@@ 4.1 Baselines
A variety of state-of-the-art LLMs and LMMs, along with the adapted forms of web-oriented models, were employed to evaluate their performance in the MMInA benchmark. The default parameters were used from either open-sourced pretrained models or API-based models for evaluation. For the web-trained agent WebShop, which was built on the specific, static environment, we utilized the GPT-3.5-turbo model to generate formatted queries to replace the phrases that should have originally been queried from the built-in environment. In text-only baselines, only textual inputs are given the LMM Fuyu-8B. We follow the display setting from [53] in a viewport $1280\times 2048$, and the webpage accessibility tree as the text input for the models. For text-based models, we performed the tasks into $2$ settings: 1) text-only: only the available textual information was given while neglecting the image information; 2) caption-augmented: using BLIP-2 [20] model to translate the information of images into text format to combine with the original text information as the input to models. For multimodal models, both the images and text information from the website were provided. We categorized the models as follows:
LLMs as Agents
Several works have shown that LLMs can act as powerful agents that have the ability to predict feasible actions and interact with the environment by prompting [25, 53, 28]. As the textual input is the accessibility tree representation of the webpage, we categorized the text-based agents into 4 groups: 1) pretrained open-source LLMs, like CodeLLaMA[37]; 2) text decoders from pretrained open-source LMM, like Fuyu-8B [4]; 3) API-based LLMs, like GPT-4 [29] and Gemini-Pro [42]; 4) pretrained web-based agents with only the language module enabled, like WebShop [49].
LMMs as Agents
With the evolution of LMMs, particularly in their remarkable capabilities for multimodal comprehension and reasoning, we embarked on a series of experiments to integrate multimodal information into task evaluation. Our experiments involved prominent LMMs such as Fuyu-8B [4], Gemini-Pro-Vision [1], and GPT-4V [2].
Heuristic-Based Web Agents
Previously, several heuristic-based web agents were specifically crafted with the intention of navigation and completion of web-based tasks [49, 8, 52]. To test the generalization ability of this category of agents, we adapted WebShop in both LLM and LMM settings with the help of GPT-3.5-turbo to format the input into the required format.
Human Baselines
We conducted a comparison of hop and task performances within the same settings as an average of three human test takers. The test takers come from various socioeconomic backgrounds, without information on the tasks before the evaluation. Human baselines consistently outperform all existing web agents with significant margins.
@@ 4.2 Main Results
The results for the different models are shown in Tab. 3, where the hop performance and task performance are evaluated respectively. The hop success rate was calculated by the percentage of successful visits for the targeted websites, and the task success rate was evaluated based on the final task completion results, i.e., the percentage of successful tasks completed by agents with the total number of tasks.
From the experimental results, we found current state-of-the-art models showed significantly degraded performance regarding the multihop tasks, which reflected their inability to effectively recognize and comprehend structured, long-context information from the web that encapsulates the fundamental details of web content. Moreover, current agents exhibited huge performance drops with the increased number of hops, revealing their limitations in long-chain reasoning capabilities.
The hop success rate, which counts every successful task completion at a website, serves as an auxiliary metric to more accurately represent the procedural performance of each agent. On single-hop tasks, GPT-4V outperformed all other models, showcasing its superior capabilities in image and context comprehension, as well as planning. However, for tasks with a hop count ranging from 2 to 4, we observed unexpected performances. Specifically, Gemini-Pro without captions and GPT-4 without captions exhibited higher performances compared to their counterparts that are augmented with captions. Further analysis of the agents’ trajectories revealed that when agents are given relatively simple tasks while being under-informed or lacking sufficient information, they tend to ”wander” through the given hops, which often results in a non-ending loop and ultimately leads to failure. This phenomenon explains why some agents exhibit a higher hop success rate while simultaneously having a low task success rate. The insight also justifies the need for a holistic evaluation protocol with MMInA.
The experimental results showed that: 1) Multimodality-reliance: Multimodal models exhibit overall higher performance in both hop and task performance that make more accurate predictions on the proposed benchmark; 2) Context window length: Language models specifically trained to comprehend structured and long contexts, like CodeLLaMA and GPT series, which are tailored for interpreting a highly structured text format, are more suitable to execute web-based tasks that utilize structured representations of webpages; 3) Web-based models: the models that trained on web-based content (e.g., Fuyu-8B, WebShop) still exhibit the versatility and adaptability in unfamiliar environments.
4.3 Analysis: Why are Multihop Web Tasks Challenging?
@@ H.C. SR(%) 1st 2nd 3rd 4th 5th 6th
Table 4: Evaluation by Hops. Tables (a) and (b) display the hop success rates (SR), distinguished by hop counts (H.C.) of the tasks ranging from 2 to 6, for the models GPT-4V and Gemini-Pro-Vision respectively. Higher success rates are marked with darker colors. Both agents fail on the early hops when solving tasks of more hops.
Search Space
Agents often exhibit poor performance on multihop tasks, struggling to complete even the initial hops. However, when each hop is isolated into a separate single-hop task, agents generally perform well on them all.
Upon analyzing the result, we discovered that for single-hop tasks, where only one reference URL is provided, agents typically explore solely within the specific website, eventually completing the task through trial and error. In contrast, for multi-hop tasks, we provide a prompt containing all the websites necessary to complete the task. When agents fail to complete a task within the expected website, they may navigate to other websites in an attempt to proceed. This excessive amount of observation can confuse the agent’s focus, leading to a catastrophic decrease in task completion rates.
For instance, in a task like ”Help me book a flight to Tokyo, find a tour guide, search for YouTube videos, rent a car, and book a hotel,” the agent may struggle with the initial step of entering the destination and clicking ”search” on a flight booking website. It might then attempt to navigate to a ”trip guide” webpage, and realize it cannot complete the task. Despite returning to the flight booking website, the agent’s limited memory prevents it from recalling previous actions in multi-step tasks. As a result, the agent may repeat actions in a cycle within the task, illustrating the challenges of multi-step tasks.
Additionally, we provide the agent with the termination condition for each hop as the prompt. However, the agent occasionally struggles to extract this crucial information, leading to a lack of awareness of the correct endpoint. As a result, it may become stuck in the current hop even after completing it, failing to proceed to the next hop to accomplish the task.
Agent Input Length
The total hop count of a multihop task defines the length of the task. It is assumed that the success rate of each hop is contingent upon the successful completion of the preceding hop, and the overall number of hops should not influence the success rate of any individual hop, provided that the hop is defined within a specific domain. However, after aligning the first hop semantically across all tasks, our empirical findings in Tab. 4 indicated an unexpected pattern that was contrary to the assumption above. We observed that agents performed better on tasks with fewer total hops, achieving higher success rates in completing the first hop. Conversely, as the total hop count increased, there was a noticeable decline in the success rate within the first hop. We attribute this phenomenon to the enlarged search space and the agent’s weak zero-shot long-context reasoning capabilities, which we resolve in Sec. 4.4.
This observation underscores the complexity inherent in multihop tasks, which extends beyond merely accumulating the performance of single-hop tasks. It poses a unique challenge that requires the coordination, handling, and control of the entire task flow. This complexity highlights the need for web agents to develop enhanced planning and reasoning capabilities to adeptly manage and execute these intricate tasks.
4.4 Memory-augmented Agents
@@
Agents operating within dynamic environments must execute actions that are informed by not only real-time environment observations and user queries but also their historical action patterns. Based on the previous experimental results, the complexity of predicting actions lies in the requirement for diverse types of memory at different phases of this process, emphasizing the importance of retaining information through various tasks, actions, and web interactions. We propose memory-augmented web agents, characterized by three specific memory systems: semantic memory, episodic memory, and procedural memory. Semantic memory comprises the agent’s general world knowledge, which can be continually sourced and updated from the Internet or comprehensive knowledge bases. In practice, semantic memory is encoded in LLM weights. Episodic memory acts as a repository to temporarily store the captured step-wise action trajectories of individual tasks, allowing agents to methodically record and subsequently recall each predictive step when predicting the next possible action for the specific task. In practice, it is usually in the form of the previous context for an autoregressive model or in-context examples. Procedural memory, activated upon the completion of a task’s action sequence, encodes the full trajectory and outcomes (success or failure) to distill historical strategies. In this work, we highlight the importance of procedural memory that significantly improves a web agent’s performance by replaying the agent’s action trajectories on past tasks similar to the current one (Tab. 3). With the three memory systems integrated, multimodal agents are equipped to retrieve a wealth of relevant information, using it as a reference point to navigate and adapt to forthcoming actions, thereby ensuring a sophisticated, context-aware approach to complex task resolution.
Implementation
We implement memory-augmented agents on top of LMMs by concatenating the action trajectories (including the task descriptions and the web contents as observations) of the last $K$ tasks to the prompt of LMMs in a sequence of tasks. The replayed experience potentially grounds the agent’s reasoning, i.e., prunes the search spaces of the agent. However, it extends the input length by $K$ times, which poses a challenge to LLMs that are trained on short corpora. To balance both sides, we explore the optimal $K$ value to construct the procedural memory. As shown in Fig. 5, the optimal performance is usually obtained at $K=2$. Our experiments revealed that agents benefiting from procedural memory references displayed superior performance in both action prediction and execution. Nevertheless, our findings also highlighted a non-linear relationship between performance and the number of historical references. In simpler tasks, such as those within the shopping or Wikipedia domain, superior performance was observed with a small number of history references, particularly when the history number was set at $1$ and $2$ respectively, where the performance with a higher history number showed limited improvements from the base conditions. With the higher history number settings, the unprioritized history will introduce biases and turbulence inside the model decision process as well, causing a lower increase in the model’s final performance. Notably, we performed experiments with Gemini-Pro-Vision, but the proposed method is model-agnostic to be flexibly deployed with any LMM or LLM.
@@ 5 Conclusion, Challenges, and Outlook
We present MMInA, a benchmark featuring three properties: 1) we benchmark agents on real-world websites and propose 1,050 multimodal multihop tasks among 14 various websites. We experiment with current SOTA LLMs and LMMs and also provide human baselines; 2) we propose a novel holistic evaluation method for multihop tasks that assess both task and hop success rates; 3) we propose a flexible memory-augmented method to improve the performance by enhancing procedural memory of agents.
Limitations
Due to the protection mechanisms employed by web pages, it’s exceptionally challenging to find a website that allows us to directly fetch images from HTML files. So one of the websites we utilized is an offline standalone website and the other is an open-source website.
@@
With the increasing prevalence of mobile devices, we aim to expand the task domain to include mobile platforms. Additionally, to enhance the capabilities of the agent, we will introduce a mechanism for long-term memory, allowing the agent to selectively remember useful actions taken during a task. For multihop tasks, our current evaluation methods primarily rely on keyword queries within website URLs or indirect approaches for assessment. Moving forward, we will consider employing an evaluation method focused on actions, which will directly guide the agent’s operations.
Potential Broader Impact
Our benchmark provides a testing bed for subsequent agent research and identifies certain issues that can guide future related studies. We do not identify any potential negative impact of this work.
Appendix A MMInA Benchmark Details
A.1 Website Links
Table A1: Links to MMInA websites. All of them are evolving real-world websites with content updated over time except Shopping. Our flexible evaluation protocol facilitates supporting more websites in the future.
  • 1
Appendix B Experiment Details
B.1 Supplementary Results
@@
We follow the previous settings of hop analysis in the main paper, illustrating the agent performance of a GPT-4V agent in Table A2. We observed again that agents performed better on tasks with fewer total hops, achieving higher success rates in completing the first hop. Conversely, as the total hop count increased, there was a noticeable decline in the success rate within the first hop. Because there are fewer long-range (¿7 hops) tasks, the success rates fluctuate due to randomness.
@@
Table A2: Performance of GPT-4V on multihop tasks based on the hop number ranging from 2 to 10. The success rate (sr) is calculated based on single-hop evaluation results over the whole completed hop numbers.
B.2 More Related Works

### v2

@@ 4 Experiments
[…]
Input Type Agent Inputs Hop Success Rate ($\uparrow$) Task Success Rate ($\uparrow$)
@@
CodeLLaMA-7B 5.71 0 0 1.61 5.71 0 0 2.79
[…]
GPT-4 38.58 20.70 3.43 13.50 38.58 3.79 0 19.85
Multimodal CogAgent-9B 6.92 0 0 1.06 6.92 0 0 3.35
[…]
@@
[…]
@@ 4.1 Baselines
We employed a variety of state-of-the-art LLMs, LMMs, and adapted web-oriented models to evaluate their performance on the MMInA benchmark. For text-based models, we conducted evaluations in two settings: 1) text-only: Only textual information from the website was used, ignoring image content. 2) caption-augmented: In addition to the text, we used the BLIP-2 Li et al. (2023a) model to generate captions for website images, incorporating visual information. For multimodal models, both image and text information from the website were provided. More details regarding the environment, model parameters, and versions are included in Sec. A. We categorized the models as follows:
1) LLM/LMM as Agent’s Reasoning Backbone
Large language / multimodal models can act as powerful backbones in agents’ reasoning processes that can predict feasible next-step actions by prompting Liu et al. (2023b); Zhou et al. (2024). As the textual input is the accessibility tree representation of the webpage, we categorized the text-based agents into 4 groups: 1) pretrained open-sourced LLMs, like CodeLLaMARoziere et al. (2023) and DeepSeek-R1 Guo et al. (2025); 2) text decoders from pretrained open-sourced LMMs, like Fuyu-8b Bavishi et al. (2023); 3) API-based LLMs, like GPT-4 Achiam et al. (2023) and Gemini-Pro Team et al. (2023). Our experiments also involved prominent LMMs such as Fuyu-8b Bavishi et al. (2023), Gemini-Pro-Vision Team et al. (2023), and GPT-4V Achiam et al. (2023).
2) Heuristic-Based Web Agents
Several heuristic-based web agents were specifically crafted with the intention of navigation and completion of web-based tasks Yao et al. (2022); Deng et al. (2023); Hong et al. (2024); Zheng et al. (2024). We selected WebShop and CogAgent as baselines to evaluate how models trained on web-based tasks perform on the MMInA benchmark.
3) Human Baselines
We conducted a comparison of hop and task performances within the same settings with an average of $3$ human test takers. The test takers come from various socioeconomic backgrounds, without information on the tasks before the evaluation. Human baselines consistently outperform all existing web agents with significant margins.
@@ 4.2 Main Results
The results for the different models are shown in Tab. 2, where the hop performance and task performance are evaluated respectively. The hop success rate reflects the percentage of successful visits to targeted websites, while the task success rate measures the percentage of tasks successfully completed by agents out of the total number of tasks.
Our experimental results indicate that current state-of-the-art models exhibit significantly reduced performance on multihop tasks. This gap highlights their struggle to effectively recognize and process structured, long-context information from web pages, which is essential for understanding web content. Additionally, performance drops sharply as the number of hops increases, revealing the models’ limitations in long-chain reasoning.
The hop success rate, which counts every successful task completion at a website, serves as an auxiliary metric to represent the procedural performance of each agent more accurately. On single-hop tasks, DeepSeek-R1-Distill-Qwen-32B outperformed all other models, indicating that the reasoning model possesses strong potential in image and context comprehension, as well as planning. However, for tasks with a hop count ranging from 2 to 4, we observed unexpected performances. Specifically, Gemini-Pro without captions and GPT-4 without captions exhibited higher performances compared to their counterparts that are augmented with captions. Further analysis of the agents’ trajectories revealed that when agents were assigned relatively simple tasks while being under-informed or lacking sufficient information, they tended to “wander” through the given hops. This often led to an endless loop, causing them to lose focus on the task’s original intent and ultimately resulting in failure. This phenomenon explains why some agents achieve a higher hop success rate while simultaneously exhibiting a low task success rate. This insight further justifies the need for a holistic evaluation protocol with MMInA.
The experimental results showed that: 1) Multimodality-reliance: Multimodal models exhibit overall higher performance in both hop and task performance, which makes more accurate predictions on the proposed benchmark; 2) Context window length: Language models like CodeLLaMA and the GPT series, designed for structured and long-context processing, excel in web-based tasks relying on structured webpage representations. Reasoning models like DeepSeek-R1 perform well in single-hop tasks due to strong contextual comprehension. However, when tackling multi-hop tasks that require retaining longer contexts, R1 struggles and exhibits degraded performance; 3) Web-based models: the models that were trained on web-based content (e.g., Fuyu-8B, WebShop) still exhibit the versatility and adaptability in unfamiliar environments.
4.3 Why are Multihop Web Tasks Challenging?
Table 3: Evaluation by Hops. Tables (a) and (b) display the hop success rates (SR), distinguished by hop counts (H.C.) of the tasks ranging from 2 to 6, for the models GPT-4V and Gemini-Pro-Vision, respectively. Higher success rates are marked with darker colors. Both agents fail on the early hops when solving tasks with more hops.
@@ H.C. SR(%) 1st 2nd 3rd 4th 5th 6th
Search Space
Agents often underperform on multihop tasks, failing at early hops, yet excel when each hop is treated as an individual single-hop task. Our analysis shows that in single-hop tasks with a single reference URL, agents tend to repeatedly attempt actions within the same website upon failure until the task is completed. Conversely, in multihop tasks where prompts include multiple websites, agents that fail on the expected site often switch to alternate sites instead of persistently retrying. This behavior leads to excessive exploration and a significant drop in task completion rates.
For example, in a complex task like “Book a flight to Tokyo, find a tour guide, watch YouTube videos, rent a car, and book a hotel”, agents may fail in the first step. Even if they correctly redirect to a relevant site, limited memory prevents them from recalling previous steps, leading to repeated actions without progress. Moreover, despite defined termination conditions for each hop, agents often fail to recognize and apply them, lingering in completed hops instead of advancing, ultimately failing the task.
Agent Input Length
The total hop count in a multihop task determines its length. Each hop’s success depends on the previous one, but the success rate of any individual hop should be independent of the total number of hops, provided it remains within the specific domain. However, after aligning the first hop semantically across all tasks, our empirical findings in Tab. 3 indicated an unexpected pattern that was contrary to the assumption above. We observed that agents performed better on tasks with fewer total hops, achieving higher success rates in completing the first hop. Conversely, as the total hop count increased, there was a noticeable decline in the success rate within the first hop. We attribute this phenomenon to the enlarged search space and the agent’s weak zero-shot long-context reasoning capabilities, which we resolve in Sec. 4.4.
These findings highlight the complexity of multihop tasks, which is not simply the sum of single-hop performances but involves intricate task flow management. This complexity underlines the necessity for web agents to possess advanced planning and reasoning skills to effectively navigate and execute multihop tasks.
4.4 Memory-augmented Agents
Agents in dynamic environments must make decisions based on real-time observations, user queries, and past trajectories. Our experiments reveal the complexity of action prediction, requiring different memory types at various stages. This highlights the need to retain information across tasks, actions, and web interactions. We propose memory-augmented web agents with three key memory systems: semantic, episodic, and procedural.
  • Semantic memory stores the agent’s general world knowledge, continuously updated from the Internet or knowledge bases, typically encoded in the weights of large language models.
  • Episodic memory temporarily holds step-by-step action trajectories, enabling the agent to recall previous actions for ongoing tasks, often represented as context in autoregressive models or in-context examples.
  • Procedural memory activates after completing a task, encoding the full action sequence and outcomes to refine strategies for future tasks.
Our results emphasize the role of procedural memory in improving agent performance by replaying past action trajectories for similar tasks (see Table 2). With these memory systems, multimodal agents can access and apply relevant information, enabling more sophisticated, contextually aware responses to dynamic environments, thereby significantly enhancing performance and adaptability.
@@ 5 Conclusion, Challenges, and Outlook
We present MMInA, a benchmark with three key features: 1) It benchmarks agents on real-world websites with 1,050 multimodal multihop tasks across 14 diverse websites, including experiments with state-of-the-art LLMs and LMMs, as well as human baselines; 2) It introduces a novel holistic evaluation method for multihop tasks, assessing both task-level and hop-level success rates; 3) It proposes a flexible memory-augmented approach to enhance agents’ performance by improving their procedural memory.
@@
Moving forward, we will consider employing an evaluation method focused on actions, which will directly guide the agent’s operations.
6 Limitations
Due to the protection mechanisms employed by web pages, it’s exceptionally challenging to find a website that allows us to directly fetch images from HTML files. So, one of the websites we utilized is an offline standalone website, and the other is an open-source website.
7 Ethical Considerations
Bias in the base multimodal models can lead to inaccurate or unfair outcomes. Users should consider the representativeness of training data to avoid biased behavior.
Appendix A Experiment Details
A.1 Environment & Parameter Settings
We followed the display settings from Zhou et al. (2024), using a viewport of $1280\times 2048$, and provided the webpage accessibility tree as text input to the models. Default parameters were used from either open-source pre-trained models or API-based models. For the web-trained agent WebShop, which was designed for a static environment and trained with specially structured queries, we used the GPT-3.5-turbo model to generate formatted queries in place of those typically sourced from the built-in environment. For the versions of API-based models, the GPT-4 model in the paper is referred to gpt-4-0125-preview; the GPT-4o model is referred to gpt-4o-2024-11-20; the GPT-4V model is referred to gpt-4-vision-preview; the Gemini-Pro model is referred to gemini-1.0-pro-001; the Gemini-Pro-Vision is referred to gemini-1.0-pro-vision-001.
A.2 Computing Resources
We used 1 Nvidia RTX6000 Ada GPU with 48 GB memory for the pre-trained baseline models (most of them have 7b/8b/9b parameters) that run the inference code locally. The inference of each epoch on MMInA takes 4-8 hours, varying by model inference performance and forward times.
A.3 Supplementary Results
@@
We follow the previous settings of hop analysis in the main paper, illustrating the agent performance of a GPT-4V agent in Tab. A1. We observed again that agents performed better on tasks with fewer total hops, achieving higher success rates in completing the first hop. Conversely, as the total hop count increased, there was a noticeable decline in the success rate within the first hop. Because there are fewer long-range (¿7 hops) tasks, the success rates fluctuate due to randomness.
Ablation of agents’ memory
We enhance LMMs with memory by appending action trajectories from the last $K$ tasks, comprising task descriptions and web content observations, to their prompts. This approach, which integrates replayed experiences, helps narrow the agent’s search space, thereby grounding its reasoning. However, this technique multiplies the input length by $K$, presenting a challenge for LMMs accustomed to shorter inputs. To find a balance, we determined the optimal $K$ value for constructing procedural memory, which, as illustrated in is typically $K=2$. Our tests show that agents with procedural memory enhancements perform better in action prediction and execution. Yet, we observed a non-linear relationship between the number of historical references and performance. In simpler tasks within domains like shopping or Wikipedia, a smaller historical set—specifically $K=1$ or $2$—tended to yield better results, while larger histories offered diminishing returns and introduced biases and disturbances into the decision-making process. Although these experiments were conducted using Gemini-Pro-Vision, our method is designed to be model-agnostic, adaptable to any LMM or LLM.
Table A1: Performance of GPT-4V on multihop tasks based on the hop number ranging from 2 to 10. The success rate (sr) is calculated based on single-hop evaluation results over the whole completed number of hops.
@@
Appendix B MMInA Benchmark Details
B.1 Datasets
B.1.1 Annotators
MMInA is constructed by three human annotators from scratch. They, varying in age and gender but proficient in web browsing, were provided pre-annotated examples and followed consistent guidelines. Each labeled different dataset portions, and cross-validation was conducted for task diversity and answer accuracy. All annotators signed formal agreements and were trained for annotation.
B.1.2 Anontation Protocals
The final trajectory includes all necessary website nodes. Agents, however, can freely explore and visit unnecessary websites before completing tasks.
B.1.3 Data Statistics
From the hop counts shown in Fig. 3, we observe that as the number of hops increases, the number of actions required by the agent also increases. However, it’s worth noting that the average number of actions required for 5-hop data is lower than that for 4-hop data. In our dataset, 4-hop content involves comparative operations, such as “Which one has a Ferris wheel in the center of the city, Tianjin or Chengdu?” Since our definition of multi-hop tasks involves navigation across different web pages, we do not categorize these comparative questions as 5-hop tasks.
Most of our multitasking revolves around travel. First, we need to determine the travel destination. We let the agent determine the travel destination by retrieving and answering questions on Wikipedia. A set of tasks is related to travel, including booking flights, reserving hotels, exchanging currencies, etc. Then we randomly select some tasks from the task set and combine them into a complete and smooth task. In each task’s JSON file, we include keywords such as “flight”, which is the three-letter code of the destination airport. The reason for these keywords is that for specific tasks, some sub-tasks cannot be measured with a unified standard, so we add these keywords to judge the endpoints of specific tasks. Another type of task related to cooking includes purchasing food on Amazon and searching for recipes. Each task is close to real life and uses real and dynamic web pages for operation. We have 1050 tasks in total and each task contains a QA pair and other supportive materials. 108 QA pairs are filtered from WebQA Chang et al. (2022). Statistical details about our dataset are shown in Fig. 2.
B.2 Tasks
B.2.1 Website Links
Tab. A2 reveals the links to each website defined in MMInA.
Table A2: Links to MMInA websites. All of them are evolving real-world websites with content updated over time, except Shopping. Our flexible evaluation protocol facilitates supporting more websites in the future.
  • 1
B.2.2 Comparision with other benchmarks
The comparison of our benchmark and the VWA benchmark is shown in A3.
Table A3: Comparison between VWA Koh et al. (2024) and MMInA. MMInA requires multimodal inputs at multiple steps to accomplish the task, which makes it a more challenging multimodal benchmark.
Websites Task
VWA
[Uncaptioned image] When did the programming language that has the largest variance in salary first appear? Answer using the information from the Wikipedia site in the second tab.
MMInA
[Uncaptioned image] Do both LIYFF-Stools Modern Home Office Chair and GGHHJ Adjustable Rotating Salon Stool PU Leather Round Rolling Stool Round Chair have armrests?
Appendix C More Related Works